# CanadaFireSat: Toward high-resolution wildfire forecasting with multiple modalities

### v1

@@ 2.1 Sample Identification
Canada being a large country, we sampled areas across all provinces and territories. As our fire labels are binary, we sampled the dataset as a series of positive and negative examples. For fire (positive) sample identification, we first extract all fires occurring between 2015 and 2023, as described in Section 2.1.1. Then, samples not including any fire event (negative) are sampled across the same period for all provinces and territories, depending on their fire weather index and acquisition dates, as detailed in Section 2.1.2.
Figure 2: Burned area in Canada in millions of hectares extracted from NBAC compared to reported values from the Canadian Interagency Forest Fire Centre (CIFFC). (a) shows the annual burned area for the entire Canada from 2016 to 2023, while (b) reports the per-region burned area for 2023 only.
@@ 2.1.1 Positive Samples
Fire samples in our CanadaFireSat dataset are identified based on the fire polygons of the National Burned Area Composite [^3] Available at https://cwfis.cfs.nrcan.gc.ca/datamart/metadata/nbac (NBAC) (Hall et al., 2020) from the Canadian National Fire Database. NBAC has been compiled annually since 1972 and integrates data from Natural Resources Canada, provincial and territorial agencies, and Parks Canada, using a rule-based approach to select the most accurate burned area perimeter delineations based on post-event satellite imagery. We focus on all fires since 2015 (the launch of the Sentinel-2 satellites), with no restriction on ignition sources or other fire metadata. In Figure 2a, we report the yearly average burned area for this period from NBAC, with 2022 reaching $1.38$ mha burned and 2023 reaching $14.01$ mha burned. This outlines the two different regimes of wildfire season for our validation (2022) and test (2023) sets compared to the average from 1972 to 2015 of $\sim 2.03$ mha. In other words, 2023 was an exceptional fire season. We compared such statistics against the average burned area reported by the Canadian Interagency Forest Fire Centre (CIFFC) for 2023: 17 mha. The difference between CIFFC and NBAC reported burned area has multiple explanations. First, the CIFFC statistics are not standardized across all territorial fire management agencies, contrary to NBAC. This is directly linked to data collection timelines, as CIFFC may provide near-real-time estimates while NBAC is compiled up to 6 months after the calendar year, leaving more room for comprehensive post-fire analysis. In Figure 2b, we also break down the regional burned area statistic for the 2023 wildfire season where the most impacted provinces and territories were Québec, Northwest Territories [^4] Natural Resources Canada, that provides NBAC data includes Nunavut fires in Northwest Territories statistics., and British Columbia. It should be noted that British Columbia is often impacted by severe wildfire seasons, and often leads Canada in terms of total burned area, but 2023 was not a record year for this province. Also, we note that the most impacted regions are those with the strongest discrepancies between reported numbers from CIFFC and NBAC.
@@
Positive samples for the CanadaFireSat dataset are extracted from the NBAC fire polygons through two aggregation processes. First, through a spatial aggregation on a $2.8\;\text{km}\times 2.8\;\text{km}$ grid over Canada, where positive samples are identified as the grid entries intersecting the fire polygons. We used a small buffer around the $2.64\;\text{km}\times 2.64\;\text{km}$ tiles to avoid any potential overlap between samples due to data imprecision. Second, a temporal aggregation is performed in two steps: 1) all fires temporally overlapping inside a grid entry are accounted as a single fire occurring from the first start date to the last end date, and 2) leveraging the 8-day temporal grid from products such as NDVI from MODIS (starting each year at the 1st of January) we aggregate all fires within a spatial grid entry occurring during the same 8-day window. This is done to build our 8-day wildfire forecasting benchmark where, for a given time-step $t$, our model should predict the probability of a fire occurring in the next 8-days, i.e. from $t$ to $t+7$ included, leveraging predictors (both satellite and environmental) from $t-\Delta t$ to $t-1$. In Figure 3(a), we showcase the spatial distribution of positive samples across Canada, for a total of $n_{pos}=88^{\prime}110$ samples before any post-processing (detailed in Section 2.2). Outside of British Columbia, most fires occur in the boreal ecosystem. This pattern is very visible across Alberta, Saskatchewan, and Manitoba, where the Great Plains in southern portions of these provinces show little to no fires compared to the boreal forest in the north.
(a) Positive sample concentration across the period 2015-2023
(b) Negative sample concentration across the period 2015-2023
@@
Figure 3: Distribution of our positive samples (containing burned area) and negative samples (following our FWI-based sampling strategy) from 2015-2023, before any post-processing.
@@ 2.1.2 Negative Samples
As we aim to build our benchmark on multi-modal inputs, including satellite image time series, we are limited in storage to densely sample Canada over the whole period from 2015 to 2023. Therefore, we sampled a negative set of size $n_{neg}=2\cdot n_{pos}$ to emulate a degree of imbalance, following other datasets on wildfire forecasting (Huot et al., 2020; Prapas et al., 2021; Kondylatos et al., 2022). We sample from the same grid defined in Section 2.1.1, $G_{y,r}$, for each year $y$ between the first and last occurring fires during that year (so beyond the wildfire season), and across all regions $r$. For a given year $y>2015$ and region $r$ we avoid locations where a fire occurred in the previous years: $\bigcup_{i=2015}^{y-1}F_{i,r}$, or locations that were already selected as negative samples in the previous years: $\bigcup_{i=2015}^{y-1}N_{i,r}$. Our negative set for a given region and year can be defined as such:
@@
where $N_{y,r}$ is the set of negative samples and $S_{y,r}$ the set of potential locations in the grid. We sample $N_{y,r}$ uniformly across levels (defined by buckets) of the fire weather index: FWI:
@@
In practice, this is done by partitioning the FWI distribution into ten buckets (or bins): $[B_{1},\dots,B_{10}]$ across the FWI quantiles $[Q_{1},\dots,Q_{9}]$ such that each bucket contains approximately $10\%$ of the observations, and uniformly sampling across those buckets for $N_{y,r}$. In practice, each bucket $B_{l}$ is defined as a subset of the FWI range:
@@
In Figure 4(a), we present the annual FWI mean for our negative sample set. We see that up to the bucket: 4 with a FWI mean of $\overline{\text{FWI}}=0.62$, most negative samples will have an FWI close to 0, as the FWI distribution of available locations for negative samples consistently presents an important peak in this range. This is representative of the FWI conditions across all regions of Canada between the first and last fires of each year. Furthermore, we show in Figure 4(b) and 4(c) that across two commonly impacted regions by wildfires, Alberta and Northwest Territories, there are strong differences in FWI bucket mean value between regions, with the delta for the top bucket reaching up to $\sim 14$ in 2021. This can be explained by the higher latitude of the Northwest Territories compared to Alberta and the presence of permafrost in their northernmost areas.
@@
Figure 4: Annual FWI mean over four buckets: $\{1,4,7,10\}$ across Canada, Alberta, and Northwest Territories.
Figure 3(b) presents the spatial distribution of the sampled negative locations across all years: it shows that, per region, the negative samples are well spread spatially, contrary to the positive samples, as we aim to represent the complete patterns of fire danger conditions. British Columbia, Alberta, Saskatchewan, and Manitoba contain the highest concentration of negative samples in certain areas due to the high concentration of fires in those regions (negative samples are sampled twice as much as positive ones). On the contrary, Nunavut, Newfoundland and Labrador, and New Brunswick are less densely sampled due to a lack of fires during the analyzed period. We select in total $n_{neg}=176,650$ negative samples. Note that some of these samples will be filtered out through the post-processing procedure described in Section 2.2. In Section 4.1, we will also introduce an adversarial sampling strategy for the negative samples to study the lower-bound performance of wildfire forecasting models in such a challenging scenario for the extreme year $2023$.
@@ 2.2 Predictors
The predictors used in CanadaFireSat fall into two categories: satellite image time series and environmental data. We begin by describing the satellite image time series, a novel component that enhances high-resolution wildfire forecasting by complementing the more commonly used environmental data. We then present the environmental data.
@@ 2.2.1 Satellite Image Time Series
Figure 5: Row 1-3 Samples of Sentinel-2 input time series for 4 locations in Canada, with only the RGB bands with rescaled intensity. Row 4 Sentinel-2 images after the fire occurred. Row 5 Fire polygons used as labels with the Sentinel-2 images post-fire.
@@
To be able to forecast wildfires at a pixel resolution of $100$ m, we need high-resolution information. However, hydrometeorological fire danger predictors cannot be found at $100$ m resolution for the entirety of Canada. Therefore, we investigate the potential of multi-spectral high-resolution satellite images as proxies for fire predictors following previous literature (Pelletier et al., 2023; Yang et al., 2021). We use the 13 bands from Sentinel-2 (S2) L1C harmonized data as proxies to several known fire predictors (NDVI, soil moisture, …). We use the L1C products as they are directly available for the whole period 2015-2023 without any need for further processing. Moreover, we extract temporal data to better estimate the impact of changes in the hydrometeorological conditions on the local ecosystem.
@@
For a given sample $x_{t}\in G_{y,r}$ (the $2.8\;\text{km}\times 2.8\;\text{km}$ grid over Canada), we downloaded all the full (no missing values) S2 images of size $2.64\;\text{km}\times 2.64\;\text{km}$ centered within each grid cell $x_{t}$ between the date $t-64$ days and $t-1$ day. We excluded images with a cloud cover above $40\%$. This represents 13 images, given the average revisit time of 5 days for S2. To avoid artifacts, we use a lossless compression: LZW, and we multiplied each band intensity by a factor $1e^{-4}$ to then rescale the values to 8-bit unsigned integers. Once all S2 images were extracted, a second filter on cloud coverage was applied, based on the S2 cloud probability product, but focusing only on the region of interest. After this filtering, the samples $x_{t}$ with less than three S2 images or covering a period of less than 40 days were removed. Finally, as multiple S2 tiles can cover a sample $x_{t}$, we kept the tile with the most valid images for this sample. The final positive sample set after filtering is of size $69,876\sim 79\%$ of the original set, and the negative sample set is of size $107,925\sim 61\%$ of the original set.
@@
The fire polygons associated with each positive sample are rasterized based on the B3 band of the S2 image that preceded the start of the fire. This process outputs binary maps of size $264\times 264$ pixels at a resolution of $10\;m$ that will then be downscale during training. Figure 5 shows examples of S2 image time series from the positive set and the expected output when a fire occurred.
@@ 2.2.2 Environmental Predictors
Fire weather indices and most wildfire forecast models rely on well-known hydrometeorological drivers such as temperature, precipitation, and soil moisture. Some forecast models also leverage vegetation indices such as NDVI, EVI, or LAI. Since we aim for high-resolution wildfire forecasting, those predictors at coarse resolutions could be seen as unnecessary compared to satellite image time series. Nonetheless, we argue that despite an important difference in resolutions between those predictors and our target outputs, multi-modal methods can benefit from coarse environmental predictors, as those are directly linked to fire probability.
For this purpose, we extract a series of environmental indices at coarse resolution, summarized in Table 2.
@@
    First, we extract five different MODIS products that describe the vegetation state and temperature at moderate-to-coarse-resolution: $500$ m and $1$ km. Vegetation indices are 8-day or 16-day composites, which similarly to SeasFire (Prapas et al., 2022), drive the temporal aggregation over 8 days of the other environmental predictors, and the NBAC burned area polygons.
@@
    We also extracted 12 hydrometeorological drivers from ERA5-Land daily (detailed in Table 2) at coarse-resolution: $11$ km, and aggregated those variables through mean, max, and min operators on the 8-day temporal grid defined by MODIS. We extend this set of predictors with three predictors: relative humidity, vapor pressure deficit, and wind speed, computed locally from ERA5 data.
@@
    Lastly, we leverage indices related to fire danger from the Copernicus Emergency Management Service (CEMS): fire weather index, also used in negative sampling, and drought code, both from the Canadian Forest Fire Weather Index. This data is the coarsest of all our environmental predictors with a resolution of $0.25$° for both latitude and longitude.
These predictors are then post-processed to set to NaN any extreme values and aligned both spatially and non-spatially with our positive and negative samples. Similar to the satellite image time series, for each sample $x_{t}$, we extract the environmental predictors from $t-64$ days to $t-1$ day. The non-spatial alignment is done via the weighted average of a given predictor over the target grid cell. The spatial alignment is done for each predictor by extracting a small window of data centered on $x_{t}$. The window size varies depending on the source resolution. We extract windows of dimension $(32,32)$ for MODIS products at $500$ m and $(16,16)$ for MODIS product at $1$km. Moreover, for ERA5-Land data, we extract windows of size $(32,32)$, and $(13,13)$ for CEMS.

### v2

@@ 2.1 Sample Identification
Covering the entirety of Canada with Sentinel-2 images at $10$ m requires extremely high storage capacity, beyond the size of typical datasets. As such, to represent all territories and provinces of Canada, we build CanadaFireSat by resorting to a sampling strategy. As our fire labels are binary, we sample the dataset as a series of positive and negative examples. For fire (positive) sample identification, we first extract all fires that occurred between 2015 and 2023, as described in Section 2.1.1. Then, samples not including any fire event (negative) are sampled across the same period for all provinces and territories, depending on their FWI and acquisition dates, as detailed in Section 2.1.2.
Figure 2: Burned area in Canada in millions of hectares extracted from NBAC, compared to the values reported by the Canadian Interagency Forest Fire Centre (CIFFC). (a) shows the annual burned area for Canada from 2016 to 2023. The difference between CIFFC and NBAC reported burned area has multiple explanations. First, the CIFFC statistics are not standardized across all territorial fire management agencies, contrary to NBAC. This is directly linked to data collection timelines, as CIFFC may provide near-real-time estimates while NBAC is compiled up to 6 months after the calendar year, leaving more room for comprehensive post-fire analysis. (b) reports the per-region burned area for 2023 only, where the most impacted provinces and territories were Québec, Northwest Territories (Natural Resources Canada, which provides NBAC data, includes Nunavut fires in Northwest Territories statistics), and British Columbia. We note that the most impacted regions are those with the strongest discrepancies between reported numbers from CIFFC and NBAC.
@@ 2.1.1 Positive Samples
Fire samples in our CanadaFireSat dataset are identified based on the fire polygons of the National Burned Area Composite [^2] Available at https://cwfis.cfs.nrcan.gc.ca/datamart/metadata/nbac (NBAC) (Hall et al., 2020) from the Canadian National Fire Database. NBAC has been compiled annually since 1972 and integrates data from Natural Resources Canada, provincial and territorial agencies, and Parks Canada, using a rule-based approach to select the most accurate data source to delineate the burned area perimeters; this includes ground and aerial surveys or post-event satellite imagery analysis from Landsat (5, 6, 7, 8, 9 or MSS), Sentinel-2, MODIS, VIIRS, and AVHRR. We focus on all fires since 2015 (the launch of the first Sentinel-2 satellite) up to 2023, with no restriction on ignition sources or other fire metadata. Over this time, a large majority of the polygons were compiled from ground survey, Landsat, aerial survey, and Sentinel-2 in this respective order. In Figure 2a, we report the NBAC yearly average burned area for this period, with 2022 reaching $1.38$ mha burned and 2023 reaching $14.01$ mha burned. This outlines the difference in wildfire season severity for our validation (2022) and test (2023) sets compared to the average from 1972 to 2015 of $\sim 2.03$ mha. In other words, 2023 was an exceptional fire season.
@@
Positive samples for the CanadaFireSat dataset are extracted from the NBAC fire polygons through two aggregation processes. First, through a spatial aggregation on a $2.8\;\text{km}\times 2.8\;\text{km}$ grid over Canada, where positive samples are identified as the grid entries intersecting the fire polygons. We used a small buffer around the $2.64\;\text{km}\times 2.64\;\text{km}$ Sentinel-2 tiles to avoid any potential overlap between samples due to imprecision in the data processing. Second, a temporal aggregation is performed in two steps: 1) all fires temporally overlapping inside a grid entry are accounted as a single fire occurring from the first fire start date to the last fire end date, and 2) leveraging the 8-day temporal grid from products such as NDVI from MODIS (starting each year at the 1st of January) we aggregate all fires within a spatial grid entry occurring during the same 8-day window. This is done to build our 8-day wildfire forecasting benchmark where, for a given time-step $t$, our model should predict the probability of a fire occurring in the next 8 days, i.e. from $t$ to $t+7$ included, leveraging predictors (both satellite and environmental, see Section 2.2) from $t-\Delta t$ to t1t-1. In Figure 3(a), we showcase the spatial distribution of positive samples across Canada, for a total of $n_{pos}=88,110$ samples before any post-processing (detailed in Section 2.2). Outside of British Columbia, most fires occur in the boreal ecosystem. This pattern is very visible across Alberta, Saskatchewan, and Manitoba, where the Great Plains in the southern portions of these provinces show little to no fires compared to the boreal forest in the north.
(a) Positive sample distribution across the period 2015-2023
(b) Negative sample distribution across the period 2015-2023
@@
Figure 3: Distribution of positive (containing burned area) and negative samples (following our FWI-based sampling strategy) from 2015-2023, before any post-processing.
@@ 2.1.2 Negative Samples
As we aim to build our benchmark on multi-modal inputs, including satellite image time series, we are limited in disk storage to densely sample Canada over the whole period from 2015 to 2023. Therefore, we sample a negative set of size $n_{neg}=2\cdot n_{pos}$ to match the degree of imbalance of other wildfire forecasting datasets (Huot et al., 2020; Prapas et al., 2021; Kondylatos et al., 2022). We sample from the same grid defined in Section 2.1.1, $G_{y,r}$, for each year $y$ between the first and last fires during that year (so beyond the wildfire season), and across all regions $r$. For a given year $y>2015$ and region $r$, we avoid locations where a fire occurred in the previous years: $\bigcup_{i=2015}^{y-1}F_{i,r}$, or locations that were already selected as negative samples in the previous years: $\bigcup_{i=2015}^{y-1}N_{i,r}$. Our negative set for a given region and year can be defined as:
@@
where $N_{y,r}$ is the set of negative samples and $S_{y,r}$ the set of potential locations in the grid. We sample $N_{y,r}$ uniformly across levels (defined by decile bins) of the FWI:
@@
In practice, this is done by partitioning the FWI distribution into ten decile bins: $B_{1},\dots,B_{10}]$ across the FWI quantiles $Q_{1},\dots,Q_{9}]$ such that each bin contains approximately $10\%$ of the observations, and uniformly sampling across those decile bins for $N_{y,r}$. Each bin $B_{l}$ is defined as a subset of the FWI range:
@@
Figure 3(b) presents the spatial distribution of the sampled negative locations across all years: it shows that, per region, the negative samples are well spread spatially, contrary to the positive samples, as we aim to represent the complete patterns of fire danger conditions. British Columbia, Alberta, Saskatchewan, and Manitoba contain the highest concentration of negative samples in certain areas due to the high concentration of fires in those regions (negative samples are sampled twice as much as positive ones). On the contrary, Nunavut, Newfoundland and Labrador, and New Brunswick are less densely sampled due to a lack of fires during the analyzed period. We select in total $n_{neg}=176,650$ negative samples that, combined with our positive samples $n_{pos}$, consitute the CanadaFireSat Train (2016 - 2022), Val (2022), and Test (2023) sets. Note that some of these samples will be filtered out through the post-processing procedure described in Section 2.2.
In Figure 4(a), we present the annual FWI mean for the negative sample set. We see that up to the decile bin number $4$ with a FWI mean: $\overline{\text{FWI}}=0.62$, most negative samples will have an FWI close to 0, as the FWI distribution of available locations for negative samples consistently presents an important peak in this range. This is representative of the FWI conditions across all regions of Canada between the first and last fires of each year. Furthermore, we show in Figure 4(b) and 4(c) that across two commonly impacted regions by wildfires (Alberta and the Northwest Territories), there are strong differences in FWI decile bin mean value between regions, with the delta for the top decile bin reaching up to $\sim 14$ in 2021. This can be explained by the higher latitude of the Northwest Territories compared to Alberta and the presence of permafrost in their northernmost areas.
@@
Figure 4: Annual FWI mean over four decile bins: $\{1,4,7,10\}$ across Canada, Alberta, and Northwest Territories.
We also observe a strong inter-annual variability between 2022 and 2023, as the latter was a record-breaking wildfire season in Canada (Jain et al., 2024), resulting in $19.6\%$ of fire patches in the Test set compared to $11.5\%$ in the Val set. This distribution shift shows that, despite similar fire weather conditions as presented in Figure 4(a), wildfire occurrence is significantly higher in the Test set compared to the Training set. This can lead to the overestimation of the performance of wildfire forecasting models: by looking at the distribution of positive and negative samples in Figure 5, one can observe that the FWI alone is a highly discriminative feature for the class fire (see Section 4.1). As a result, we introduce an adversarial sampling strategy for the negative samples to study the lower-bound performance of wildfire forecasting models for the extreme year $2023$, named Test Hard. In this adversarial test set, we aim to make the distribution of the negative population similar to that of the positive population with respect to the FWI, making ignition the main discriminative factor. To sample negative samples for Test Hard, we perform a stratified sampling for the year 2023 in the following way. First, by extending Equation 1 to account for both land cover and the month of the year. Then, for a given land cover $c$ and month of the year $m$, we sample $N_{y,r,m,c}$ uniformly across levels (defined by decile bins) of the FWI for the positive samples population $F_{y,r,m,c}$ :
$P_{FWI}(x|N_{y,r,m,c})\propto P_{FWI}(x|F_{y,r,m,c})$ (4)
and sample $n_{neg}(y,r,m,c)\simeq 2\times n_{pos}(y,r,m,c)$ negatives. The land cover is downloaded from ESA WorldCover at $10$ m for 2020. The resulting distribution is shown in Figure 5 and represents 77,247 complementary negative samples to CanadaFireSat statistics reported in Table 1. By deploying the trained networks on Test Hard, where ignition acts as the main distriminative factor, we can assess their performance on modeling this complex triggering factor whose patterns can only be implicitly learned from the training data. For this reason, the performance of our trained models for high-resolution wildfire forecasting on Test Hard can be considered as a lower bound for such an extreme wildfire season as presented in Section 4.1. Further details about the distribution of samples across land-cover classes are provided in Figure 17.
@@ 2.2 Predictors
The predictors used in CanadaFireSat fall into two categories: satellite image time series and environmental data.
@@ 2.2.1 Satellite Image Time Series
Figure 6: Row 1-3 Samples of Sentinel-2 input time series for four locations in Canada. We show only the RGB bands at $10$ m resolution with rescaled intensity. Row 4 Sentinel-2 images after the fire occurred. Row 5 Fire polygons used as labels with the Sentinel-2 images post-fire as background.
@@
To be able to forecast wildfires at a patch resolution of $100$ m, we need high-resolution information. However, hydrometeorological fire danger predictors cannot be found at $100$ m resolution for the entirety of Canada. Therefore, we investigate the potential of multi-spectral high-resolution satellite images as proxies for fire predictors following previous literature (Pelletier et al., 2023; Yang et al., 2021). We use the 13 bands from Sentinel-2 (S2) L1C harmonized data as proxies to several known fire predictors, such as NDVI or soil moisture. We use the L1C products as they are directly available for the whole period 2015-2023 without any need for further processing. Moreover, we extract temporal data to better estimate the impact of changes in the hydrometeorological conditions on the local ecosystem. Top-of-atmosphere reflectance from Sentinel-2 is impacted by aerosols, clouds, topography effects, and other phenomena that can bias its measurement across the multi-spectral bands for numerous land cover types (Sola et al., 2018), in particular for shorter wavelengths like the RGB bands. This can impact the computation of radiometric indices often used in burned area mapping (Howe et al., 2022) and the precision of non-local and multi-temporal analyses of Sentinel-2 data. However, machine and deep learning can largely mitigate those limitations by implicitly learning the approximate corrections necessary for the downstream application targeted through correction agnostic models (Rußwurm et al., 2023; Wright et al., 2025) or L1C specific ones (Medina-Lopez, 2020; Wright et al., 2024) even in the context of burned area mapping (Rumora et al., 2020). We hypothesize that our models can mitigate the lack of atmospheric corrections for the task of wildfire forecasting.
@@
For a given sample $x_{t}\in G_{y,r}$ (the $2.8\;\text{km}\times 2.8\;\text{km}$ grid over Canada), we download all the full (no missing values) S2 images of size $2.64\;\text{km}\times 2.64\;\text{km}$ following (Manas et al., 2021) centered within each grid cell $x_{t}$ between the date t64t-64 days and t1t-1 day. We exclude images with a cloud cover above $40\%$. This represents 13 images, given the average revisit time of 5 days for S2 (after the launch of Sentinel-2B). To avoid artifacts, we use a lossless compression, and we multiplied each band intensity by a factor $1e^{-4}$ to then rescale the values to 8-bit unsigned integers. Once all S2 images are extracted, a second filter on cloud coverage was applied, based on the S2 cloud probability product, but focusing only on the sample location. After this filtering, the samples $x_{t}$ with less than three S2 images or covering a period of less than 40 days are removed, as we aim to learn local temporal dynamics. Finally, as multiple S2 tiles can cover a sample $x_{t}$, we keep the tile with the most valid images for this sample. The final positive sample set after filtering is of size $69,876$ ($\sim 79\%$ of the original set), and the negative sample set is of size $107,925$ ($\sim 61\%$ of the original set). For Test Hard, the final number of negative samples is $66,406$ ($\sim 86\%$ of the previously identified samples).
@@
The fire polygons associated with each positive sample are rasterized based on the B3 band of the S2 image that preceded the start of the fire. This process outputs binary maps of size $264\times 264$ pixels at a resolution of $10$ m that will then be downscaled to $100$ m resolution during training. Figure 6 shows examples of S2 image time series from the positive set and the expected output when a fire occurred (last row).
@@ 2.2.2 Environmental Predictors
Fire weather indices and most wildfire forecast models rely on hydrometeorological drivers such as temperature, precipitation, and soil moisture. Some forecast models also leverage vegetation indices such as NDVI, EVI, or LAI. We include such coarse environmental predictors (summarized in Table 2) despite the difference in resolution between them and our target outputs, since we believe that multi-modal methods can benefit from them, as they are strongly correlated to fire probability.
@@
    First, we extract five different MODIS products from MOD15A2H, MOD11A1, and MOD13A1 that describe the vegetation state and temperature at moderate-to-coarse-resolution: $500$ m and $1$ km. Vegetation indices are 8-day or 16-day composites, which, similarly to SeasFire (Prapas et al., 2022), drive the temporal aggregation over 8 days of the other environmental predictors, and the NBAC burned area polygons.
@@
    We also extracted 12 hydrometeorological drivers from ERA5-Land daily (detailed in Table 2) at coarse-resolution ($11$ km), and aggregated those variables through mean, max, and min operators on the 8-day temporal grid defined by MODIS. We extend this set of predictors with three additional ones: relative humidity, vapor pressure deficit, and wind speed, computed locally from ERA5 data.
@@
    Lastly, we leverage indices related to fire danger from the Copernicus Emergency Management Service (CEMS): FWI, also used in negative sampling, and drought code, both from the Canadian Forest Fire Weather Index. This data is the coarsest of all our environmental predictors with a resolution of 0.250.25° for both latitude and longitude.
These predictors are then post-processed to set to NaN any extreme values and aligned both spatially and non-spatially with our positive and negative samples. Similar to the satellite image time series, for each sample $x_{t}$, we extract the environmental predictors from t64t-64 days to t1t-1. The non-spatial alignment is done via the weighted average of a given predictor over the target grid cell. The spatial alignment is done for each predictor by extracting a small window of data centered on $x_{t}$. The window size varies depending on the source resolution. We extract windows of dimension $(32,32)$ for MODIS products at $500$ m and $(16,16)$ for MODIS product at $1$ km. Moreover, for ERA5-Land data, we extract windows of size $(32,32)$, and $(13,13)$ for CEMS. As a consequence, for a given sample $x_{t}$ CanadaFireSat provides spatial predictors at multiple scales covering different spatial contexts. Models trained on CanadaFireSat should consider this difference in scale across modalities, as those presented in Section 3.